OpenStudy (anonymous):

Im trying to derive the following... ((4x+2)^2 - 4) / (x+1) And I use quotient rule to get...

OpenStudy (anonymous):

\[\frac{ 8(4x+2)(x+1) - (4x+2)^2 -4 }{ (x+1)^2 }\]

OpenStudy (anonymous):

But now I need help with simplification

OpenStudy (welshfella):

you are correct so far

OpenStudy (welshfella):

well first expand those expressions in parentheses

OpenStudy (ybarrap):

Be sure you multiply out the numerator first and then simplify that. You will it cancel x+1 in the denominator and left with something simple.

OpenStudy (campbell_st):

well then factor the numerator (4x + 2){8(x + 1) -(4x +2)+4)} simplify the part inside { }

OpenStudy (anonymous):

Oh alright

OpenStudy (anonymous):

the inside would be 4x+10 right?

OpenStudy (campbell_st):

that's what I got

OpenStudy (anonymous):

then do i expand those 2 brackets?

OpenStudy (campbell_st):

well it really depends... as the factored numerator is a correct solution... I'd leave it that way to avoid errors... but it really depends on what the question as for

OpenStudy (anonymous):

it just asks for the derivative and apparently the answer is 16 but i can't get that

OpenStudy (welshfella):

16?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

well there must be a condition and you have been given a value of x to substitute the question may ask for the slope of the curve at x = -2 or something like that

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

Nope none of that

OpenStudy (welshfella):

i think its 16 if x = 1 - if my simplification is correct

OpenStudy (campbell_st):

are you asked to find f '(1) or anything like that

OpenStudy (welshfella):

f'(x) is a function of x

OpenStudy (anonymous):

No xD lol

OpenStudy (anonymous):

do you think it may be wrong?

OpenStudy (welshfella):

then the answer cant be 16

OpenStudy (welshfella):

yes - something is wrong...

OpenStudy (campbell_st):

well if you are asked for the derivative you have it correct. if you are asked to find the slope at x = 1 or f'(1) then you would need a numeric answer

OpenStudy (lynfran):

u forget to take the derivative of the denominator that the problem here

OpenStudy (anonymous):

Ohh i get it! I have to simplify the numerator before even differentiating it! Cause that way I get f(x) = 16x and f'(x) = 16

OpenStudy (campbell_st):

the derivative of the denominator is 1

OpenStudy (welshfella):

yes thats right

OpenStudy (anonymous):

\[f(x) = \frac{ 16x^2+16x+4-4 }{ (x+1) } = \frac{ 16x(x+1) }{ (x+1) } = 16x\]

OpenStudy (ybarrap):

Perfect! $$ f(x)=\cfrac{(4x+2)^2 - 4}{x+1}=\cfrac{16x^2+16x+4-4}{x+1}=\cfrac{16x(x+1)}{x+1}\\ $$ Now you see it: \(f'(x)=16\)

OpenStudy (anonymous):

Yeah exactly!! Thanks so much for the help! :D

OpenStudy (welshfella):

oh yes!!!! i missed that

OpenStudy (welshfella):

what a crafty question!

OpenStudy (anonymous):

Lol i know right! XD

OpenStudy (welshfella):

well spotted

OpenStudy (welshfella):

that function is a 'disguised' straight line with a slope of 16

OpenStudy (anonymous):

Ahahaha xD yep!

OpenStudy (welshfella):

you should be able to get 16 using the quotient rule but i didn't . There must have been an error in my work.